OpenStudy (lanadelbey):

Find S15 for the series -1 + -3 + -5 + -7 +...

OpenStudy (sooobored):

The sum? or the 15th term in the sequence?

OpenStudy (dumbcow):

Sum of an arithmetic series: \[S_n = \frac{n(a_1 + a_n)}{2}\] where \[a_n = a_1 + d(n-1)\] \[a_1 = -1\] \[d = -2\]